OpenStudy (mony01):

1. Which of the graphs correctly displays the positions of the motorcycle and car as functions of time? 2. How long does it take from the moment when the motorcycle starts to accelerate until it catches up with the car? In other words, find t2-t1. 3. How far does the motorcycle travel from the moment it starts to accelerate (at time t1) until it catches up with the car (at time t2)?

OpenStudy (mony01):

A motorcycle is following a car that is traveling at constant speed on a straight highway. Initially, the car and the motorcycle are both traveling at the same speed of 23.0m/s , and the distance between them is 50.0m . After t1 = 3.00s , the motorcycle starts to accelerate at a rate of 6.00m/s2 . The motorcycle catches up with the car at some time t2.

OpenStudy (ahsome):

1) Graph A

OpenStudy (ahsome):

2) T1=3 Seconds, T2=6 Seconds Time difference+ \[T1-T2\] \[6-3=3\]

OpenStudy (ahsome):

3)\[Distance=Speed*Time\] \[D=6*3\] \[D=18m\]

OpenStudy (ahsome):

Do you understand @mony01?

OpenStudy (mony01):

hold on for number 2 so you're saying that T2 is 6?

OpenStudy (mony01):

I don't think thats right because the answer you got was wrong

OpenStudy (ahsome):

Ok then. That means we need to find T2, not assuming that T2 is T1*2

OpenStudy (mony01):

how we do that?

OpenStudy (ahsome):

You know that he the motorbike closes the gap between the car at 6.00ms^2 acceleration

OpenStudy (ahsome):

The gap is 50 metres

OpenStudy (mony01):

so we divide 6 by 50?

OpenStudy (ahsome):

@hartnn We need your help

hartnn (hartnn):

s = ut +1/2 at^2 s=50 u= 23 a=6 find t that will be t2

OpenStudy (mony01):

how did you get this equation?

OpenStudy (ahsome):

This equation is one of the basic Physics equation

OpenStudy (mony01):

Can someone verify if T2 is 50/3?

OpenStudy (ahsome):

Don't you have the answers with you?

OpenStudy (mony01):

no

OpenStudy (mony01):

@hartnn

OpenStudy (ahsome):

s = ut +1/2 at^2 s=50 u= 23 a=6 \[s=ut+\frac{ 1 }{ 2 }at^2\] \[50=23t+\frac{ 1 }{ 2 }6t^2\] \[50=23t+3t^2\] \[3t^2+23t-50=0\]

OpenStudy (mony01):

thats the part i got stuck on how i solve for t?

OpenStudy (ahsome):

Use the quadratic equation \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (ahsome):

\[3t^2+23t-50\] The A value is 3, B value is 23, C value is -50

OpenStudy (ahsome):

Substitute into the equation: \[\frac{-b\pm\sqrt{b^2-4ac}} {2a}\] \[\frac{-23\pm\sqrt{23^2-4*3*-50}} {2*3}\]

OpenStudy (ahsome):

Do you understand @mony01?

OpenStudy (mony01):

yea i got 1.767 is it right?

OpenStudy (ahsome):

I haven't done the sum myself, but If you did it right, then yeah :)

OpenStudy (mony01):

do i still subtract 3

OpenStudy (ahsome):

Yes, you have to

OpenStudy (ahsome):

Remember, this equation has 2 answers.

OpenStudy (mony01):

yea one is negative do i use it?

OpenStudy (ahsome):

No, never use negative number. Use positive

OpenStudy (ahsome):

Only use negative when it isn't real-life, like in a graph

OpenStudy (mony01):

@Ahsome thanks I found a solution without having to use the quadratic formula thank you anyways

OpenStudy (ahsome):

Oh, Good job :D (How did you do it?)